Quantrex Quantrex AcademyJEE · NEET · NDA PYQs with solutions Open app
NDA 2025
Exam
NDA
Year
2025
Subject
Mathematics
Chapter
Probability

NDA 2025 · Mathematics · Probability

The probability of a man hitting a target is (1)/(5). If the man...

The probability of a man hitting a target is 15.Ifthemanfires7times,thenwhatistheprobabilitythathehitsthetargetatleasttwice?\dfrac{1}{5}. If the man fires 7 times, then what is the probability that he hits the target at least twice?

Options

  1. A1−(35)(45)61 - \left(\dfrac{3}{5}\right)\left(\dfrac{4}{5}\right)^6
  2. B1−(35)(45)71 - \left(\dfrac{3}{5}\right)\left(\dfrac{4}{5}\right)^7
  3. C1−(115)(45)61 - \left(\dfrac{11}{5}\right)\left(\dfrac{4}{5}\right)^6
  4. D1−(115)(45)71 - \left(\dfrac{11}{5}\right)\left(\dfrac{4}{5}\right)^7

Correct answer

C. 1−(115)(45)61 - \left(\dfrac{11}{5}\right)\left(\dfrac{4}{5}\right)^6

Step-by-step solution

Let pp be the probability of hitting the target and qq be the probability of missing the target.

p=15p = \dfrac{1}{5}

q=1−p=45q = 1 - p = \dfrac{4}{5}

Number of trials, n=7n = 7

Let XX be the random variable representing the number of hits. The probability of hitting the target at least twice is given by:

P(X≥2)=1−P(X<2)P(X \ge 2) = 1 - P(X \lt 2)

P(X≥2)=1−[P(X=0)+P(X=1)]P(X \ge 2) = 1 - [P(X = 0) + P(X = 1)]

Using the binomial distribution formula P(X=r)=nCrprqn−rP(X = r) = ^{n}C_{r} p^r q^{n-r}:

P(X=0)=7C0(15)0(45)7=(45)7P(X = 0) = ^{7}C_{0} \left(\dfrac{1}{5}\right)^0 \left(\dfrac{4}{5}\right)^7 = \left(\dfrac{4}{5}\right)^7

P(X=1)=7C1(15)1(45)6=75(45)6P(X = 1) = ^{7}C_{1} \left(\dfrac{1}{5}\right)^1 \left(\dfrac{4}{5}\right)^6 = \dfrac{7}{5} \left(\dfrac{4}{5}\right)^6

Substituting these values:

P(X≥2)=1−[(45)7+75(45)6]<br><br>P(X \ge 2) = 1 - \left[ \left(\dfrac{4}{5}\right)^7 + \dfrac{7}{5} \left(\dfrac{4}{5}\right)^6 \right]<br><br>P(X ≥ 2) = 1 - \left(\dfrac{4}{5}\right)^6[45+75]6 \left[ \dfrac{4}{5} + \dfrac{7}{5} \right]

P(X≥2)=1−(115)(45)6P(X \ge 2) = 1 - \left(\dfrac{11}{5}\right) \left(\dfrac{4}{5}\right)^6

Answer: 1−(115)(45)61 - \left(\dfrac{11}{5}\right)\left(\dfrac{4}{5}\right)^6
Practice Probability on Quantrex Academy →

More from Probability

Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026Three events A , B and C are such that A and B are disjoint, A and C are independent, B… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026For a Binomial distribution with mean 6 and standard deviation 2 , what is the value of… NDA 2026Passage: A 4-digit number is selected at random formed by using the digits 0 , 1 , 2 , 3… NDA 2026In an entrance test there are multiple choice questions. There are four options for each… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Consider the following statements in respect of the events A , B , C : I. (A B C) ( A B C… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026The probability that a family owns a laptop is 0.68 ; that it also owns a desktop is 0.56… NDA 2025An urn contains 10 white and 5 red balls. If two balls are drawn at random, then what is… NDA 2025An urn contains 5 white, 6 red and 4 blue balls. Three balls are drawn at random. What is… NDA 2025Under which of the following conditions may binomial distribution be used? I. The number… NDA 2025Three faces of a die are black, two faces are white and one face is red. The die is… NDA 2025A person X speaks the truth 4 out of 5 times and person Y speaks the truth 5 out of 6… NDA 2025 Full Probability list Probability JEE Maths PYQsJEE Mathematics PYQs All NDA PYQs